What are the X-ray and -y components of vector Ď. The magnitude of the vector is D=3m, and the angle α=45 degree b) what are the X- and Y-axis components of vector Ē the magnitude of the vector is E-4.5m, and the angle
β=37 degree
1
Expert's answer
2019-07-10T10:39:54-0400

(a)

"D_x=D\cos\alpha=3.00\cos 45^{\circ}=2.12""D_y=D\sin\alpha=3.00\sin 45^{\circ}=2.12"

(b)

"E_x=E\cos\beta=4.50\cos 37^{\circ}=3.60""E_y=E\sin\beta=4.50\sin 37^{\circ}=2.71"